### v2

# MCGrad: Multicalibration at Web Scale

Abstract.
We propose MCGrad, a novel and scalable multicalibration algorithm. Multicalibration - calibration in sub-groups of the data - is an important property for the performance of machine learning-based systems. Existing multicalibration methods have thus far received limited traction in industry. We argue that this is because existing methods (1) require such subgroups to be manually specified, which ML practitioners often struggle with, (2) are not scalable, or (3) may harm other notions of model performance such as log loss and Area Under the Precision-Recall Curve (PRAUC). MCGrad does not require explicit specification of protected groups, is scalable, and often improves other ML evaluation metrics instead of harming them. MCGrad has been in production at Meta, and is now part of hundreds of production models. We present results from these deployments as well as results on public datasets.
Multicalibration, Calibration, Uncertainty Quantification, Fairness
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
A machine learning model is said to be calibrated when its predictions match the true outcome frequencies (Silva Filho et al., 2023; Kull et al., 2017; Bella et al., 2010). The importance of calibration to ML-based systems has been widely recognized (Kendall and Gal, 2017; Guo et al., 2017; Kull et al., 2019), including in web applications such as content ranking (Kweon et al., 2022; Penha and Hauff, 2021), recommender systems (Gulsoy et al., 2025; da Silva and Jannach, 2025; Steck, 2018), digital advertising (Chen and Qi, 2022; Chaudhuri et al., 2017; Fan et al., 2023) and content moderation (Liu et al., 2025; Vidgen et al., 2020; Kivlichan et al., 2021). Calibration is necessary for such systems to make optimal decisions without under- or overestimating risk or opportunity.
Multicalibration is a more powerful property that extends the concept of calibration to ensure that predictors are simultaneously calibrated on various (potentially overlapping) groups (Hebert-Johnson et al., 2018; Pfisterer et al., 2021; Noarov and Roth, 2023; Jung et al., 2021). Often, multicalibration is applied as a post-processing step to fix a base classifier’s predictions. While it originated in the study of algorithmic fairness, its significance extends far beyond it. A growing body of work shows that multicalibration can increase model performance and robustness in many circumstances, ranging from out-of-distribution prediction to confidence scoring of LLMs  (Gopalan et al., 2022b; Yang et al., 2024; Wu et al., 2024; Kim et al., 2022; Detommaso et al., 2024; Dwork et al., 2022).
Despite its potential, multicalibration has received limited traction in industry. We argue that this is due to three main reasons. First, existing multicalibration methods require protected groups to be manually defined (Jin et al., 2025): not only do users need to specify the covariates that require protection (e.g., user age and user country), but also they must define concrete protected groups through binary membership indicators, such as “is an adult in the US?”. This poses a significant challenge to real-world application of multicalibration because practitioners (a) may not have established the precise set of protected groups, (b) may be subject changes to the definition of protected groups over time thus requiring ongoing effort and change management (e.g., due to changes related to legal context, policy frameworks, ethical considerations, or otherwise), or (c) may only seek an overall performance improvement without prioritizing specific groups for protection. The difficulties defining protected groups become even more pronounced when the set of features that we want to protect is large. Second, existing methods lack the ability to scale to large datasets and/or many protected groups, which makes them hard to deploy in production. For instance, Hebert-Johnson et al. (2018)’s algorithm scales at least linearly in time and memory with the number of groups, being potentially inefficient when a large set of groups is specified. To be deployed at web scale, multicalibration methods must be highly optimized so as to not introduce significant computational requirements at training or inference time. Third, existing multicalibration methods lack guardrails to be safely deployed, and often introduce a risk that multicalibrating a predictor ends up harming model performance may prevent deployment. This may be caused by overfitting.
In this paper, we introduce MCGrad (MultiCalibration Gradient Boosting) a novel multicalibration algorithm deployed in production that is lightweight and safe to use. MCGrad requires only the specification of a set of protected features, rather than predefined protected groups. It then identifies miscalibrated regions within this feature space, and calibrates a base classifier’s predictions over all groups that can be defined based on these features. On a high level, MCGrad uses a Gradient Boosted Decision Tree (GBDT) algorithm recursively, such that the result converges to a multicalibrated predictor (Section 3.1). By using a highly optimized GBDT implementation, like LightGBM (Ke et al., 2017), MCGrad inherits scalability and regularization (Section 3.2). Finally, it employs an early stopping procedure that guarantees, with high probability, that it improves the base predictor’s likelihood and does not overfit, ensuring that model performance is not harmed (Section 3.3).
Empirically, on benchmark datasets, we show that MCGrad improves the base predictor’s outputs in terms of multicalibration, calibration, and even predictive performance (Section 4). We complement this analysis by describing the impact of MCGrad on real-world data: MCGrad has been used in production at Meta on hundreds of machine learning models, and in total generates over a million multicalibrated real-time predictions per second (Section 5). To the best of our knowledge, this is the large-scale adoption of multicalibration in production. With these real-world results, we add to the growing body of evidence that shows that multicalibration can significantly contribute to ML model performance. We also provide valuable practical learnings from applying multicalibration in industry (Section 6), and link multicalibration to related areas (Section 7).
2. Background on (Multi)Calibration
Let $\mathcal{X}=\mathbb{R}^{d}$ be a $d$-dimensional input space and $\mathcal{Y}=\{0,1\}$ the binary target space. Let $D=\{(x_{i},y_{i})\}_{i=1}^{n}\sim p(X,Y)$ be the dataset with $n$ i.i.d. samples, where the random variables $X$, $Y$ represent, respectively, a $d$-dimensional feature vector and the target. A probabilistic predictor $f$ is a map $f\colon\mathcal{X}\to[0,1]$, which assigns an instance $x\sim p(X)$ to an estimate of its class conditional probability $p(Y=1\mid X=x)$. For any probabilistic predictor $f$, we denote with $F$ its logit (referred to as the predictor), i.e., $f(x)=\mu(F(x))$, where $\mu(F)=1/(1+e^{-F})$ is the sigmoid function. We assume a base probabilistic predictor $f_{0}$ is given; this predictor is the one targeted for calibration via a post-processing procedure.
Let $\mathcal{L}\colon\mathbb{R}\times\mathcal{Y}\to\mathbb{R}$ be the log loss (negative log-likelihood), i.e., for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and any predictor $F$,
$\mathcal{L}(F(x),y)=-[y\log(\mu(F(x)))+(1-y)\log(1-\mu(F(x)))].$
We use the symbol $\mathbb{E}$ for the expectation with respect to an arbitrary distribution that is clear from the context, and $\mathbb{E}_{D}$ for the sample mean over $D$.
Calibration
A predictor $f$ is perfectly calibrated if and only if $\mathbb{P}(Y=1\mid f(X)=p)=p$, where $p$ is the true underlying probability. Intuitively, for all input pairs $(x,y)\in D$, if $f$ predicts $0.8$, we expect that $80\%$ of them have $1$ as label. Recently, Tygert (2023) introduced the Estimated Cumulative Calibration Error (ECCE), a novel parameter-free metric for measuring calibration that computes the maximum difference between the means of labels and predictions over any score interval:
(1) $\textsc{ECCE}(f)=\frac{1}{n}\max_{1\leq i\leq j\leq n}\big|\sum_{k=i}^{j}(y_{(k)}-f(x_{(k)}))\big|,$
where $(x_{(1)},y_{(i)}),\ldots,(x_{(n)},y_{(n)})$ are the points $\{(x_{i},y_{i})\}_{i\in D}$ ordered such that $f(x_{(1)})\leq\cdots\leq f(x_{(n)})$.
This metric enjoys useful statistical properties. In particular, it can be standardized by dividing it by the following scaling statistic: [^1] Notice that $\sigma(f)$ is the standard deviation of the sample mean of prediction labels with prediction probabilities according to $f$. For a calibrated predictor $f$, the expected value of $\nicefrac{{\text{ECCE}(f)}}{{\sigma(f)}}$ scales as $c$, for large data size $n$, with $c=2\sqrt{2/\pi}\approx 1.6$ (Arrieta-Ibarra et al., 2022).
$\sigma(f)=\sqrt{\frac{1}{n^{2}}\sum_{k=1}^{n}f(x_{k})(1-f(x_{k}))}.$
This gives rise to two ways of interpreting the metric. The absolute scale $\text{ECCE}(f)$ is designed to measure the magnitude of miscalibration. When the data size increases, $n\rightarrow\infty$, $\text{ECCE}(f)$ converges to a true magnitude of miscalibration (0 if $f$ is calibrated and a positive value otherwise). The sigma scale $\nicefrac{{\text{ECCE}(f)}}{{\sigma(f)}}$ however converges to a universal distribution when $f$ is calibrated, and diverges to infinity when it is miscalibrated. The two scales are used to answer different questions. ECCE is used to answer the question ”is the miscalibration large?”. $\nicefrac{{\text{ECCE}(f)}}{{\sigma(f)}}$ is used to answer the question ”is there statistical evidence of miscalibration?”. For example, we might have that $\text{ECCE}(f_{1})=0.05=20\sigma(f_{1})$ and $\text{ECCE}(f_{2})=0.05=2\sigma(f_{2})$. This would tell us that while the magnitude of miscalibration on the data appears to be similar for $f_{1},f_{2}$, the result is statistically significant for $f_{1}$ but not for $f_{2}$ could simply be due to random sampling. See (Arrieta-Ibarra et al., 2022; Tygert, 2023) for further discussion.
Multicalibration
Unlike calibration, multicalibration considers the intersection of the score interval and the protected groups. Let $\mathcal{G}\subset\{g:[0,1]\rightarrow\{0,1\}\}$ be the set of interval membership functions, i.e. $g(v)=\mathbbm{1}_{I}(v)$ for $v\in[0,1]$, and any interval $I\subseteq[0,1]$. Similarly, let $\mathcal{H}\subseteq\{h:\mathcal{X}\rightarrow\{0,1\}\}$ be a set of group membership functions over $\mathcal{X}$, such that a point $x\in\mathcal{X}$ belongs to a group $h\in\mathcal{H}$ if and only if $h(x)=1$. The groups can be overlapping, i.e., some point $x\in\mathcal{X}$ may belong to multiple groups.
Multicalibration is defined in various ways across the literature (Hebert-Johnson et al., 2018; Shabat et al., 2020; Błasiok et al., 2024; Gopalan et al., 2022b; Deng et al., 2023; Haghtalab et al., 2023a). Here, we employ an operational definition of multicalibration:
Definition 2.1 (Multicalibration).
A probabilistic predictor $f$ is $\alpha$-multicalibrated ($\alpha$-MC) with respect to $\mathcal{H}$ if, for the MC-deviation given as $\Delta_{h,g}(f)\coloneqq\big|\mathbb{E}\left[h(X)g(f(X))(Y-f(X))\right]\big|$ and a given scale parameter $\tau_{h}(f)$:
$\Delta_{h,g}(f)\leq\alpha\tau_{h}(f)\quad\forall h\in\mathcal{H},g\in\mathcal{G}.$
We say that $f$ is multicalibrated if $\alpha=0$.
A common choice is to use a uniform bound on the MC-deviation, which is accommodated by choosing the scale parameter $\tau_{h}(f)$ to be a constant equal to $1$. In Appendix A.1 we show that Definition 2.1, with $\tau_{h}(f)^{2}=\mathbb{E}\left[h(X)f(X)(1-f(X))\right]$, naturally matches with Multicalibration Error (MCE) by Guy et al. (2025), which quantifies multicalibration error by computing the maximum ECCE over the protected groups with a suitable normalization:
(2) $\text{MCE}(f)=\max_{h\in\mathcal{H}}\frac{\text{ECCE}_{h}(f)}{\sigma_{h}(f)}$
where $\text{ECCE}_{h}(f),\sigma_{h}(f)$ are defined as $\text{ECCE}(f),\sigma(f)$ above but restricted to the group $h$. Roughly speaking, MCE measures the strongest statistical evidence of miscalibration across protected groups. The unit of the MCE is the same as for $\text{ECCE}/\sigma$. It can be re-scaled to an absolute measure by multiplying with $\sigma(f)$.
Using a multicalibration metric that is equivalent to the definition of $\alpha$-MC can help ensure that empirical work is aligned with theory. In Section 4 we use MCE to measure multicalibration, thereby directly estimating the minimal $\alpha$ for which a model is $\alpha$-MC. To our knowledge, this is the first empirical multicalibration study that achieves this. See Appendix A.1 for further discussion.
3. MCGrad: A Practical Algorithm for Multicalibration
This paper tackles the following problem:
Given:
a labeled dataset $D$, an initial probabilistic predictor $f_{0}$;
Return:
a probabilistic predictor $f$ that is $\alpha$-MC w.r.t. $\mathcal{H}$.
Designing an algorithm for multicalibration without specifying a set of groups is challenging for four reasons. First, the absence of groups requires the multicalibration model to automatically determine the regions of the feature space where the base model is highly miscalibrated. Second, adjusting the model’s output to improve its calibration in a specific region might harm the predictions over other regions, thus increasing the multicalibration error. Third, a post-processing algorithm has to be fast and lightweight to prevent large memory consumption or increased latency of model inference. Fourth, while existing algorithms claim to achieve multicalibration on the training set, they are subject to overfitting issues.
a probabilistic predictor $f_{0}$, a dataset $D$;
$D_{train}$, $D_{valid}$ = train_validation_split($D$)
$F_{0}=\mu^{-1}(f_{0})$ # inverse sigmoid transformation
$\varepsilon_{-1}=+\infty$, $\varepsilon_{0}=\mathbb{E}_{D_{valid}}\left[\mathcal{L}(F_{0}(X),Y)\right]$ # initialize errors
$t=1$ # initialize number of rounds for early stopping
while $\varepsilon_{t-1}-\varepsilon_{t}>0$ do
  $h_{t}(x,f_{t-1}(x))=$ fit GBDT on $\{((x_{i},f_{t-1}(x_{i})),y_{i})\}_{i\in D_{train}}$
  $\theta_{t}=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{D_{train}}\left[\mathcal{L}(\theta\cdot(F_{t-1}(X)+h_{t}(X,f_{t-1}(X))),Y)\right]$
  $F_{t}(x)=\theta_{t}\cdot(F_{t-1}(x)+h_{t}(x,f_{t-1}(x)))$
  $f_{t}(x)=\mu(F_{t}(x))$
  $\varepsilon_{t+1}=\mathbb{E}_{D_{valid}}\left[\mathcal{L}(F_{t}(X),Y)\right]$
  $t\leftarrow t+1$
end while
#\# Found t1t-1 to be the best number of rounds
for $s$ in $1,\dots,t-1]$ do
  $h_{s}(x,f_{s-1})=$ fit GBDT on $\{((x_{i},f_{s-1}(x_{i})),y_{i})\}_{i\in D}$
  $\theta_{s}=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{D}\left[\mathcal{L}(\theta\cdot(F_{s-1}(X)+h_{s}(X,f_{s-1}(X))),Y)\right]$
  $F_{s}(x)=\theta_{s}\cdot(F_{s-1}(x)+h_{s}(x,f_{s-1}(x)))$
  $f_{s}(x)=\mu(F_{s}(x))$
end for
$f(x)=\mu(F_{t-1}(x))$
$f$;
Algorithm 1 MCGrad
Our proposed algorithm MCGrad, presented in Algorithm 1, is a lightweight algorithm that multicalibrates a base model without requiring the specification of the groups on which the model has to be calibrated. MCGrad relies on a simple key observation: by including the base model’s predictions as a feature, the loss function of the multicalibration algorithm captures the feature values that correspond to miscalibrated initial predictions. In addition, decreasing the loss function implies correcting the base model’s predictions for some regions of the feature space. Because corrections in some regions may negatively impact others, MCGrad runs multiple rounds, using the previous round’s processed predictions as input. Since GBDTs are regularized using shrinkage (or a step size <1<1), a simple rescaling step reduces the number of trees and rounds required, while having a negligible impact on overfitting. Finally, MCGrad employs early stopping to avoid overfitting.
3.1. Achieving Multicalibration with $T$ Rounds
Our key intuition builds on the following two key insights. First, GBDT returns a solution $f$ that approximately sets the gradient of the loss function $\mathcal{L}$ to zero:
$\mathbb{E}_{D}\left[h(X)(Y-f(X))\right]=0\qquad\forall h\in\mathcal{H},$
where $\mathcal{H}$ is the set of all regression trees over the feature space.
Second, if we augment the feature space $\mathcal{X}$ with a single additional feature $f_{0}(x)$, then the GBDT approximately achieves
(3) $\mathbb{E}_{D}\left[h(X,f_{0}(X))(Y-f_{1}(X))\right]=0\qquad\forall h\in\mathcal{H}^{\prime}$
where $\mathcal{H}^{\prime}\subset\{h:\mathcal{X}\times[0,1]\rightarrow\mathbb{R}\}$ is the space of trees with an extra real-valued input, and $f_{1}(x)=\mu(F_{0}(x)+h_{1}(x,f_{0}(x)))$. This creates a bridge between (a) group and interval membership functions and (b) regression trees over the augmented feature space. Strictly speaking, for any $h\in\mathcal{H},g\in\mathcal{G}$ there exists $h^{\prime}\in\mathcal{H}^{\prime}$ such that $h(x)g(f(x))=h^{\prime}(x,f(x))$ for $x\in\mathcal{X}$.
Perhaps surprisingly, Eq. (3) is equivalent to 0-multicalibration according to Definition 2.1 under the condition $f_{1}\equiv f_{0}$. However, this condition may not always hold: correcting $f_{0}$ for some regions identified by $h(x,f_{0}(x))$ returns a predictor $f_{1}$ that might be miscalibrated on regions identified by its own predictions $h(x,f_{1}(x))$. To resolve this, we create a loop over multiple rounds: in each round $t$, we train a GBDT model $f_{t}$ using the features $x$ and the previous round’s predictions $f_{t-1}$, such that, after $T$ rounds, we have
$\mathbb{E}_{D}\left[h(X,f_{T-1}(X))(Y-f_{T}(X))\right]=0\qquad\forall h\in\mathcal{H}^{*}$
where $f_{T}(x)=\mu(F_{T-1}(x)+h_{T}(x,f_{T-1}(x)))$.
Intuitively, if for some probabilistic predictor $f^{*}$, $f_{T}\rightarrow f^{*}$ as $T\rightarrow\infty$, then for all $h\in\mathcal{H}^{*}$, $\mathbb{E}_{D}\left[h(X,f_{T-1}(X))(Y-f_{T}(X))\right]\rightarrow\mathbb{E}_{D}\left[h(X,f^{*}(X))(Y-f^{*}(X))\right]=0$ as $T\rightarrow\infty$, which is our desired condition. This suggests that, for sufficiently large $T$, the probabilistic predictor $f_{T}$ should be approximately multicalibrated. See Section A.2 for further intuitions and theoretical results.
3.2. Fast Training and Prediction
To be deployed at web scale there are strict requirements for an algorithm to be efficient both at training and inference. Several design choices ensure that MCGrad meets these requirements.
Efficient gradient boosting
First, the algorithm has been designed to rely on a relatively small number of calls to a GBDT, delegating the most compute intensive steps to one of many highly optimized GBDT implementations (e.g.  (Ke et al., 2017; Chen and Guestrin, 2016)). This differentiates it from existing implementations such as (Hebert-Johnson et al., 2018). In our implentation, we use LightGBM.
Rescaling the logits.
GBDTs are regularized in multiple ways, including using a step size smaller than $1$ for scaling each additional tree. While this is helps to avoid overfitting, it results in a predictor which can be strictly improved by rescaling it by a factor slightly greater than $1$. As a result, the next round will attempt to apply this rescaling. Since linear rescaling is not easily expressed by decision trees, this may require many trees, ending up requiring unnecessarily many trees to achieve multicalibration.
We introduce a simple rescaling after every round:
$\theta_{t}=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{D}\left[\mathcal{L}(\theta\cdot(F_{t-1}(X)+h_{t}(X,f_{t-1}(X))),Y)\right]$
where $h_{t}$ is obtained with LightGBM. We call the round $t$’s learned predictor (on the logits) as $F_{t}=\theta_{t}(F_{t-1}+h_{t})$.
Note that this constant is typically very close to $1$ and has a very limited detrimental effect on regularization. Rescaling is not the same as using a step size of $1$: it affects the whole sum of trees (i.e., the whole predictor), while the step size only targets each tree sequentially (i.e., each weak learner).
3.3. Preventing Overfitting
Modern GBDT algorithms support various methods for regularization, such as limiting the growth of the trees, the leaf splits, and the number of trees. MCGrad’s recursive structure can give rise to additional overfitting beyond standard GBDTs, which we address in the following ways.
Early stopping
While the multiple rounds are required for convergence, they also increase the capacity of the model. With $T$ rounds and $M$ trees in each round, the MCGrad model is at least as expressive as a tree ensemble with $T\cdot M$ trees. Since the model capacity of tree ensembles is $\Omega(\text{number of trees})$ (see (Shalev-Shwartz and Ben-David, 2013)), overfitting is more likely for a large number of rounds.
We solve this problem with a standard early stopping procedure on the number of rounds $T$. Specifically, we split i.i.d. the dataset $D$ into training ($D_{train}$) and validation ($D_{valid}$) sets, and determine the optimal number of rounds $T$ by taking the last round before the expected loss over the validation set increases. That is,
$T=\min\{t\colon\mathbb{E}_{D_{valid}}\left[\mathcal{L}(F_{t+1}(X),Y)-\mathcal{L}(F_{t}(X),Y)\right]>0\}-1,$
where the models $f_{t}$ are obtained using $D_{train}$ instead of $D$. Using early stopping in a real-world deployed system has a relevant consequence: with high probability, MCGrad does not harm the base model’s prediction, and, instead, would select $T=0$ as optimal number of rounds if the first step decreased the initial performance:
$\mathbb{E}_{D_{valid}}\left[\mathcal{L}(F_{1}(X),Y)-\mathcal{L}(F_{0}(X),Y)\right]>0\implies T=0$
which means that $f_{T}=f_{0}$.
Regularizing through the min sum Hessian in leaf
Augmenting the data with the previous round’s model necessarily gives rise to regions of the augmented feature space that are particularly prone to overfitting. Consider a leaf of a partially constructed tree in GBDT. The leaf splitting algorithm can choose to split the leaf on values of the previous predictor $f_{t}(x)$. In this leaf, the left tail $f_{t}(x)<a$ for some small $a$ will contain only negative labels even if the true distribution may assign a positive probability to a positive label (and analogously for the right tail). In that case the new model can improve the likelihood simply by assigning a very low probability to this tail, as low as zero. Common regularization strategies, like setting the minimum number of samples per leaf, are insufficient to address this scenario. We solve this problem by using a more targeted form of regularization that limits the minimum total Hessian in a leaf, which is offered as one of multiple regularization techniques in LightGBM. The Hessian in a leaf $S$ equals $\sum_{(x,y)\in S}f_{t}(x)(1-f_{t}(x))$ and, as such, is a refined version of the simple sample size rule $|S|$. As the predicted probabilities become close to 0 or $1$, the total Hessian becomes smaller and split in multiple leaves can no longer be considered. As shown in our ablation studies (Section 4) this regularization reduces overfitting and improves performance.
4. Benchmark Experimental Analysis
We evaluate MCGrad on both public and production data. In this section, we present experiments on benchmark datasets and compare MCGrad against state-of-the-art baselines. In Section 5 we present results of MCGrad in production. Here, we focus on four research questions:
  • Q1.
    How does MCGrad compare to existing baselines on unspecified groups?
  • Q2.
    Existing methods require manual specification of protected groups. Does MCGrad still protect those specified groups, even though they are not specified in MCGrad?
  • Q3.
    Does MCGrad benefit from running multiple rounds?
  • Q4.
    What is the effect of rescaling the logits and the regularization through min sum Hessian in leaf on MCGrad?
4.1. Experimental Setup
Data.
We use a total of $11$ tabular benchmark datasets. Five of the datasets are used in  Hansen et al. (2024)’s evaluation of multicalibration methods. We add datasets by using alternative definitions of the ground truth, using other available columns in the American Community Survey (ACS) data. We describe the datasets in more detail in Appendix B. The datasets range from sample size 11k Medical Expenditure Panel Survey (MEPS) to 3.2M (ACSEmploy), and from 10 features (ACSIncome) to 139 (MEPS).
Baselines.
We compare MCGrad against two multicalibration baselines: Discretization-Free MultiCalibration (DFMC) (Jin et al., 2025) and HKRR Hebert-Johnson et al. (2018). DFMC shares some similarities with MCGrad: It fits a single GBDT with a fixed maximum depth of 2 for each weak learner. Protected groups are explicitly specified by the user, rather than learned like in MCGrad. Protected groups are provided to the algorithm as binary features that encode group membership. HKRR is the boosting procedure proposed by Hebert-Johnson et al. (2018), for which we use  Hansen et al. (2024)’s implementation [^2] https://github.com/dutchhansen/empirical-multicalibration. We also include Logistic Regression as BasePred, the base predictor $f_{0}$ shared by all methods, and Isotonic (Zadrozny and Elkan, 2002) (Isotonic Regression), a widely adopted calibration algorithm.
Metrics.
We evaluate MCGrad relative to baselines on three dimensions: the effect on 1) predictive performance, 2) calibration, and 3) multicalibration.
1) Predictive performance. We quantify using Area Under the Precision-Recall curve (PRAUC) (Davis and Goadrich, 2006) and log loss.
2) Calibration. Various calibration metrics have been proposed in the literature. The Expected Calibration Error (ECE) (Naeini et al., 2015) is the most commonly used calibration measure in machine learning, which bins the model scores and computes the deviation from perfect calibration, i.e., the absolute difference between the model’s predicted accuracy and its empirical accuracy (Naeini et al., 2015). Binning-based calibration metrics, including ECE, vary significantly based on the choice of bins. This is a well-documented shortcoming (Arrieta-Ibarra et al., 2022; Roelofs et al., 2022). Alternative metrics that aim to circumvent the drawbacks of binning include smooth ECE (smECE) which relies on kernel smoothing (Błasiok and Nakkiran, 2023), but merely replaces sensitivity to the arbitrary choice of binning with sensitivity to an arbitrary choice of kernel (Arrieta-Ibarra et al., 2022). The Estimated Cumulative Calibration Error (ECCE) (Arrieta-Ibarra et al., 2022) is a calibration metric that is rooted in cumulative statistics that does not require making an arbitrary choice in bin or kernel, which we covered in more detail in Section 2. Brier score is a proper scoring rule that is commonly used to evaluate probabilistic forecasts that is essentially just the mean squared error of predicted probability. Brier score decomposes into calibration and the model’s ability to separate the positive and negative class (Blattenberger and Lad, 1985). In this paper, we quantify calibration error using ECCE, with exceptions in Section 5, where for pragmatic reasons we resort to ECE and Brier score whenever ECCE was not measured in the production system.
3) Multicalibration. Błasiok et al. (2024) proposed the maximum group-wise smECE metric, which was used for empirical evaluations in Hansen et al. (2024). Guy et al. (2025) propose an extension of ECCE to quantify multicalibration, called Multicalibration Error (MCE), which we defined in Eq. (2). Maximum group-wise smECE and MCE have in common that they both are defined as a maximum of some quantity over groups, but differ in the calibration quantity calculated in each group, where they respectively use smECE and ECCE. In this paper, we use MCE to quantify multicalibration. In Appendix B we additionally report group-wise smECE, thereby providing results that are in-line with (Hansen et al., 2024). We can measure any multicalibration metric with respect to either of two sets of groups: Prespecified Groups: refers to a small (less than $15$ across all datasets) set of protected groups that are formed using various features of the datasets. For this, we use the groups specifications that were used in  Hansen et al. (2024). Unspecified Groups: covers the case where the user does not specify which groups to protect against. In that case, the scores are expected to be multicalibrated with respect to all possible groups, and to calculate this, an extensive set of combinatorially generated groups are used.
Hyperparameters.
MCGrad uses the same hyperparameters as in production deployment (see Appendix C). For HKRR we follow Hansen et al. (2024) and pick the best hyperparameter from a set of four specifications using a held-out validation set. For DFMC we use the default hyperparameters of LightGBM since the paper does not suggest any specific hyperparameters.
4.2. Experimental Results
Figure 1. Multicalibration Error computed using unspecified groups (left) or the manually prespecified groups (right) for all compared methods on each benchmark dataset. Overall, MCGrad achieves a better (lower) error for $10$ out of $11$ datasets when tested on unspecified groups, and for 55 out of $11$ datasets when tested on prespecified groups.
Figure 2. Improvement (in %) of MCGrad’s variants relative to the original version. While setting $T=1$ yields a significant drop in performance, the effect of the rescaling factor and min sum Hessian in leaf are mild due to the datasets limited size.
Baseline MCE log loss PRAUC ECCE
Avg Rank Avg Rank Avg Rank Avg Rank
MCGrad 6.60 1.18 0.375 1.00 0.714 1.00 1.97 2.27
Isotonic 24.00 2.91 0.410 3.27 0.662 3.91 1.53 1.54
DFMC 24.95 3.00 0.408 2.27 0.671 2.18 2.20 2.91
HKRR 25.74 3.73 0.414 4.00 0.637 5.00 3.29 3.54
BasePred 29.07 4.18 0.414 4.45 0.668 2.91 4.66 4.73
Table 1. Average value (Avg) and rank (Rank) over the datasets, for various metrics. MCGrad has the best average log loss and PRAUC, and second-best average ECCE.
Q1. MCGrad vs existing baselines on unspecified groups.
Figure 1 (left-hand side) shows the Multicalibration Error (MCE) for all compared methods using all available features. Overall, MCGrad outperforms all baselines on $10$ out of the $11$ datasets by reducing the base predictor’s MCE between $4\%$ (Credit) and $93\%$ (ACSEmploy), and obtaining an average reduction of $56.1\%$. In contrast, DFMC, Isotonic, and HKRR only achieve an average reduction of, respectively, $11.9\%$, $9.6\%$, and $1.4\%$. In addition, MCGrad never harms the base predictor’s performance: even on MEPS, where there is low evidence of BasePred’s miscalibration (3.9 MCE), it slightly improves the MCE by $2.8\%$. On the other hand, Isotonic, DFMC, and HKRR increase (worsen) the MCE of BasePred on $3$, 33 and $4$ datasets respectively.
Table 1 reports the average performance over the datasets. MCGrad achieves by far the best average log loss improvement ($10.4\%$) and PRAUC improvement ($8.1\%$) over the base predictor. Besides these average improvements over datasets MCGrad never harms performance which helps build the necessary trust among ML engineers to delpoy it in production environments. Other methods don’t have this guarantee. For instance while, DFMC improves model performance metrics by a lower $1.8\%$ and $0.76\%$, it harms PRAUC and log loss on two datasets. Isotonic and HKRR even decrease the PRAUC on average by, respectively, $1\%$ and $5\%$ with a negligible change in log loss. For each metric, we rank all compared methods from $1$ (best) to $5$ (worst) and report the average ranking position in Table 1. A lower average rank indicates how often the method is preferred, as opposed to the average performance, which reflects the magnitude. MCGrad consistently ranks as #1\#1 baseline on log loss and PRAUC, while only ranking third for MCE on the MEPS dataset. Finally, MCGrad ranks second best method in terms of global calibration (ECCE), losing against Isotonic Regression, which is expected because it is trained to directly optimize the global calibration of the base predictor.
Q2. MCGrad vs existing baselines on prespecified groups.
To verify that good performance on the MCE metric for all groups implies also protection for prespecified groups, we report results just for these groups in Figure 1 (right-hand side). Overall, MCGrad outperforms all baselines on $5$ out of $11$ datasets achieving a $54\%$ average MCE improvement, and is followed by DFMC that is the best baseline on three datasets (ACSMobility, HDMA, ACSTravel) and achieves an average MCE improvement of $51\%$. HKRR obtains an average improvement in MCE of $36\%$, while Isotonic only achieves an improvement of $16\%$. Note that HKRR and DFMC have direct access to these groups, while MCGrad only has access to the features used to define the groups.
Q3. The impact of multiple rounds on MCGrad
Because MCGrad recurrently improves its own predictions by using $T$ rounds, one relevant question is whether $T>1$ is really necessary in practice. For this goal, we compare MCGrad with its variant Only One Round that forces the number of rounds T=1T=1.333Early stopping can possibly return $T=0$. Figure 2 (purple line) shows the relative improvement of Only One Round for the four considered metrics and all benchmark datasets. In the majority of cases, it is evident that Only One Round deteriorates MCGrad’s performance: on the MCE and ECCE axes, it more than doubles its value (relative improvement <100<-100) for, respectively, $6$ and $5$ datasets, while, on the log loss and PRAUC axes, it worsens the performance by approximately $2\%$ or more on, respectively, $8$ and $6$ out of $11$ datasets. Interestingly, the largest deterioration of performance corresponds to MCGrad running for the highest number of rounds: the optimal $T$ for ACSEmploy, ACSIncome, ACSPublic, and ACSMobility is, respectively, $20$, $35$, $27$ and $22$.
Q4. Impact of rescaling the logits and the regularization through min sum Hessian in leaf
Rescaling the logits of all predictions by learning a constant speeds up MCGrad training by reducing rounds, especially on very large datasets. As a downside, it might affect MCGrad’s performance. Figure 2 (light blue bar) shows the relative deterioration after removing rescaling. Effects are mild and balanced: removal worsens MCE, log loss, and PRAUC on $6$, 66, and $7$ datasets respectively, with average drops of $9\%$, $0.5\%$, and $0.03\%$.
LightGBM’s default min sum Hessian leaf (MSHL) is $0.001$ with a min leaf sample size of $20$. We claim that higher MHL values prevent probabilities from pushing to extremes, reducing overfitting (Section 3.3). We compare MCGrad with MHL=0.001=0.001 versus our default of $20$. Figure 2 (dark blue bar) shows mild deterioration using $0.001$, likely due to limited dataset sizes. On average, lower MHL slightly harms performance, but this becomes evident on ACSEmploy (the largest dataset), where the MHL variant runs only $2$ rounds (vs. $10$) because our regularization mechanism stops the training earlier to avoid overfitting.
5. Results in Production Deployment
We describe the results obtained by using MCGrad in a real-world production system. MCGrad has been deployed and used in production at Meta. Specifically, we have integrated MCGrad into two ML training platforms at Meta: the publicly released Looper (Markov et al., 2022), and an internal platform that we refer to as MLplatform#2. The evaluation consists of a total of $147$ binary classification models that serve live production traffic: $27$ models on Looper and $120$ models on the other platform. On both ML training platforms, we run MCGrad with default hyperparameters (see Appendix C).
Looper Results.
Looper (Markov et al., 2022) is a system that automates model training, online A/B test comparisons between the new model variant and the existing production model, and launch decisions on whether or not to promote the new model variant. These launch decisions are made based on whether the new model variant outperforms the existing production model on the selected evaluation metric with statistical significance.
For 27 binary classification models that serve active production traffic on Looper, we created a variant of the model that applies MCGrad as calibration post-processing. We started online A/B tests to compare these 27 models against the active production models, all of which are GBDTs. Those production models all had previously already applied Platt scaling (Platt et al., 1999).
We found that on $24$ out of $27$ models, the variant with MCGrad statistically significantly outperformed the same model with Platt scaling (Platt et al., 1999), resulting in the promotion of the MCGrad-calibrated model variant to become the primary production model.
Looper also monitors area under the precision-recall curve (PRAUC) on online production data. PRAUC improved on $24$ of the $27$ models and was neutral on the rest. Due to data retention, we are able to calculate the exact PRAUC for four of the $27$ Looper models for which we most recently deployed MCGrad, as well as the Brier score (Brier, 1950). Table 2 summarizes the results on those models.
Model PRAUC ( $\uparrow$ ) Brier Score ( $\downarrow$ )
Platt Scaling MCGrad Platt scaling MCGrad
#1 0.5295 0.6565 (+23.9%) Not available Not available
#2 0.1824 0.2433 (+33.4%) 0.0117 0.0116 (-0.45%)
#3 0.2161 0.2300 (+6.4%) 0.0642 0.0632 (-1.56%)
#4 0.6268 0.6354 (+1.4%) 0.0394 0.0365 (-7.24%)
Table 2. Online results for four models on Looper, where either Platt Scaling or MCGrad is used for post-processing.
MLplatform#2 Results.
First Time Period Second Time Period
Percentile Logloss PRAUC AUROC ECE Logloss PRAUC AUROC ECE
10th -26.50% +3.58% +3.15% -93.88% -16.34% +6.31% +1.97% -90.84%
25th -10.35% +0.93% +0.52% -87.45% -3.43% +1.92% +1.10% -83.07%
50th -1.39% +0.32% +0.17% -53.83% -0.86% +0.06% +0.16% -58.19%
75th -0.51% +0.00% +0.03% -14.54% -0.11% -0.95% 0.0% -21.96%
90th +0.02% -2.07% -0.20% +7.56% +1.75% -4.85% -0.39% +26.23%
Table 3. Summary statistics for the two time periods of the impact of MCGrad on $120$ (first period, left) and $65$ (second period, right) ML models on MLplatform#2.
This ML training platform is widely used at Meta. ML model training scripts on this platform define the training set, a validation set, the test set, the ML model architecture, and the set of model evaluation metrics. The ML platform allows running an evaluation flow that trains the model on the training set and evaluates the model on the test set using the specified set of evaluation metrics, or an publish flow that prepares an API endpoint that can generate predictions for live production traffic. We integrated MCGrad into the evaluation flow so that we obtain an additional set of evaluation results for a variant of the model that is calibrated using MCGrad.
For every ML model evaluation on this ML platform, we collected parallel evaluation results with and without MCGrad. We conducted data collection in two separate periods. Table 3 summarizes the results. In the first data collection period, we obtained results on $120$ binary classification production models. We found that applying MCGrad post-processing improved logloss for $88.7\%$ of the models, compared to the production models without it. PRAUC improved for $76.7\%$ of the models, AUROC for $80.2\%$, and Expected Calibration Error (ECE) (Naeini et al., 2015) for $86\%$. Note that on most models where no improvement was found, there was no metric degradation either, due to the early stopping. In Table 3 we can see that that the worst 10th-percentile metric effects of MCGrad still result in degradations. We manually investigated a sample of those models, and found that these were mostly explained by errors by ML practitioners in the train and test set specification. E.g., in some cases the train and test set were clearly from a different population, or sample weights were defined in the train set but not in the test set.
In a second data collection period, we obtained results on $65$ binary classification models. We found that MCGrad provided improvements over the production models on the logloss metric for $80.3\%$ of the models, on the PRAUC metric for $75.4\%$ of the models, on the AUROC for $79.3\%$, and ECE for $86\%$.
6. Learnings from Production Deployment
Despite the recent focus of the academic community on multicalibration, it has seen little to no uptake in industry. Here we summarize various learnings from applying multicalibration in industry through production deployments of MCGrad and of the Multicalibration Error metric (MCE) (Guy et al., 2025).
Learning 1. Multicalibration has business value.
Our results demonstrate that multicalibration post-processing can improve real-world models’ multicalibration as well as performance metrics that tend to correlate with business outcomes, such as PRAUC. Furthermore, baselines such as Isotonic Regression do not achieve similar results. This learning contradicts some observations by Hansen et al. (2024): (1) Isotonic Regression is often competitive compared to algorithms designed specifically for multicalibration, and that (2) practitioners may face a trade-off between multicalibration and predictive performance. MCGrad achieves better results than baselines and rarely degrades performance.
Learning 2. Practitioners struggle to define protected groups.
Many multicalibration algorithms (Hebert-Johnson et al., 2018; Jin et al., 2025) require manual specification of the protected groups. In our experience deploying multicalibration in production, we found that many ML engineers struggle to manually define relevant protected groups, and lack guidelines and frameworks for defining them. This experience is in line with findings from a series of interviews on AI fairness adoption in industry (Holstein et al., 2019), where it turns out that practitioners often lack knowledge on different types of biases and cannot identify the relevant biases to correct for. Hence, our experience is that in order to succeed in obtaining industry adoption, a multicalibration method needs to relieve users of the burden of manually defining protected groups, like MCGrad does.
Learning 3. Practitioners consider adopting multicalibration methods only when they work out-of-the-box, reliably, and with no risk of harming model performance
Many multicalibration algorithms have hyperparameters. We found that industry practitioners are open to applying multicalibration on their production ML models, but only if the required effort is limited. Hyperparameter tuning, for example, is often seen as too complex or time-consuming. Furthermore, they perceive it as a reliability risk: if a method requires hyperparameter tuning to work well, then they wonder if the hyperparameter configuration will still work well in subsequent model training runs.
Therefore, for a multicalibration algorithm to find successful adoption in industry, it is important that it works out-of-the-box, without the need for hyperparameter tuning. Moreover, while practitioners care about multicalibration, we found that they are not willing to accept a degradation in metrics like the log loss and the PRAUC, which are often believed to relate to topline metrics.
We defined a set of hyperparameter default values for MCGrad that achieves (1) no degradation in PRAUC and log loss, and (2) substantial reduction in MCE. Appendix C details the procedure used to obtain these hyperparameters. Results in Sections 4 and 5 confirm that those default values work sufficiently well to consistently produce log loss and PRAUC improvements over models calibrated with Platt scaling (on Looper) or uncalibrated models (on MLplatform#2), and consistently yield log loss, PRAUC, and MCE improvements on the public datasets.
7. Related Work
Multicalibration. Multicalibration as a learning objective was first introduced by (Hebert-Johnson et al., 2018), requiring a predictor to remain calibrated even when conditioned on membership in any of a potentially large collection of subgroups of the domain and the predicted value.
Considerable work has been devoted to developing the theoretical foundations of multicalibration, focusing on computational complexity and elucidating its relation to other learning concepts (Detommaso et al., 2024; Gopalan et al., 2023; Deng et al., 2023; Błasiok et al., 2024).
Low-degree multicalibration (Gopalan et al., 2022b) generalizes conditioning on the prediction by using a weight function of the predicted value belonging to a class of weight functions. It was shown that lower computational complexity can be achieved by restricting the class of weight functions to low-degree polynomial functions.
Swap multicalibration (Gopalan et al., 2023) is a stronger concept that implies multicalibration. This stronger condition has been shown to be satisfied by some existing multicalibration algorithms, such as the algorithm of (Hebert-Johnson et al., 2018). Swap multicalibration has been shown to be equivalent to notions of swap agnostic learning and swap omniprediction.
Omnipredictors, a notion originally introduced in (Gopalan et al., 2022a), strengthen the classical notion of agnostic learning by requiring a predictor to provide a minimization guarantee for each loss function in a class of loss functions. This concept was extended in (Gopalan et al., 2023) to the notion of swap omnipredictors. (Gopalan et al., 2023) has shown equivalence between swap multicalibrated predictors and swap omnipredictors for the class of convex functions satisfying certain conditions.
In (Haghtalab et al., 2023b), a connection between multicalibration and multi-objective learning is noted and exploited to formulate the learning as a game dynamics to achieve state-of-the-art guarantees for a diverse set of multicalibration learning problems. Recent work (Okoroafor et al., 2025) studied omniprediction algorithms and their theoretical guarantees.
Algorithms for multicalibration. The early proposals of multicalibration algorithms, such as HKRR (Hebert-Johnson et al., 2018) (e.g., implemented in MCBoost (Pfisterer et al., 2021) or by Hansen et al. (2024)) and LSBoost (Globus-Harris et al., 2023), rely on discretizing the output space and calibrating the prediction for each discretization level until convergence. A discretization-free multicalibration algorithm was proposed in (Jin et al., 2025), requiring group functions as input and using a gradient boosting machine to fit an ensemble of depth-two decision trees to the residuals of an uncalibrated predictor using both the original features and the predictions of the uncalibrated predictor. For guaranteeing multicalibration within an additive error, this algorithm requires a loss saturation condition to hold, i.e., that the squared-error loss can be further reduced only by a small amount. Our algorithm uses gradient boosting in a similar manner, but does not require pre-specified groups, and leverages a multi-round recursive gradient boosting and allows for arbitrary user-defined decision trees to be used in each round.
8. Conclusion
We tackle the problem of designing a multicalibration algorithm without having access to pre-specified protected sets, that is fast and safe to deploy in production. These three aspects are relevant for practitioners as they are often, (1) unable to define good groups for their task, (2) forced to limit the overhead in terms of memory and speed, and (3) required to guarantee that the algorithm does not harm the base predictor. We proposed MCGrad, a novel algorithm that achieves multicalibration by optimally finding the regions of the feature space where the miscalibration is large, and fixing the predictions. Because this might create new regions with evidence of miscalibration, MCGrad employs multiple rounds, where in each round it corrects its own output. In addition, MCGrad’s implementation is fast, lightweight, and safe to deploy. We showed that MCGrad outperforms existing baselines on benchmark datasets and provided results obtained in a large-scale industry production deployment at Meta consisting of hundreds of production models. Finally, we included a set of practical takeaways from our experience applying multicalibration in industry.
Appendix A Theoretical Results
A.1. On the Multicalibration Error
Different definitions of approximate multicalibration have been considered, referred to as $\alpha$-multicalibration. One common choice is using a uniform bound on the absolute deviation, $\Delta_{h,g}(f)\leq\alpha$ where the precise choice of $h,g$ varies (Błasiok et al., 2024; Deng et al., 2023; Gopalan et al., 2022b; Haghtalab et al., 2023a). Another common choice is a uniform bound on the conditional deviation, $\delta_{h,g}(f):=|\mathbb{E}\left[h(X)g(f(X))(Y-f(X))\mid h(X)g(f(X))=1\right]|\leq\alpha$, (Hebert-Johnson et al., 2018), (Shabat et al., 2020). These definitions were improved on by (Haghtalab et al., 2023a) which showed that a tighter bound can be achieved with no additional requirements (”for free”) by defining $\alpha$-multicalibration as $\Delta_{h,g}(f)\leq\alpha\sqrt{\mathbb{E}\left[h(X)\right]}$.
We next show that Definition 2.1, for the expected deviation corresponding to the sample mean deviation over dataset $D$, i.e.,
$\mathbb{E}_{D}\left[h(X)g(f(X))(Y-f(X))\right]=\frac{1}{n}\sum_{k=1}^{n}h(x_{k})g(f(x_{k}))(y_{k}-f(x_{k})),$
and the scale parameter chosen as
(4) $\tau_{h}(f)=\sqrt{\mathbb{E}_{D}\left[h(X)f(X)(1-f(X))\right]},$
matches with the MCE metric introduced in Guy et al. (2025), defined in Equation (2).
Proposition A.0.
Given a set $\mathcal{H}$ of group membership functions, a probabilistic predictor $f$ is $\alpha$-multicalibrated with respect to $\mathcal{H}$, with the scale parameter $\tau_{h}(f)$ as in (4), if and only if $\text{MCE}(f)\leq\alpha\sqrt{n}$.
Proof.
First, note that
$\Delta_{h,g}(f)=\frac{1}{n}\big|\sum_{k=1}^{n}h(x_{k})g(f(x_{k}))(y_{k}-f(x_{k}))\big|.$
Combining with the definition of $\text{ECCE}_{h}(f)$, we get that
$\max_{g\in\mathcal{G}}\Delta_{h,g}(f)=\frac{n_{h}}{n}\text{ECCE}_{h}(f).$
Plugging in the definition of $\text{MCE}(f)$, given in Equation (2), we have
$\frac{1}{\sqrt{n}}\text{MCE}(f)$ $=$ $\frac{1}{\sqrt{n}}\max_{h\in\mathcal{H}}\frac{\text{ECCE}_{h}(f)}{\sigma_{h}(f)}$
$=$ $\max_{h\in\mathcal{H},g\in\mathcal{G}}\frac{\sqrt{n}\Delta_{h,g}(f)}{n_{h}\sigma_{h}(f)}$
$=$ $\max_{h\in\mathcal{H},g\in\mathcal{G}}\frac{\Delta_{h,g}(f)}{\tau_{h}(f)}$
Now, from the last expression, $\alpha$-MC with respect to $\mathcal{H}$ holds if and only if $\text{MCE}(f)\leq\alpha\sqrt{n}$. ∎
A.2. Convergence Analysis
In this section, we analyze the convergence of Algorithm 1 under certain conditions for the GBM used in each round of the algorithm. Specifically, we show two results. First, we show conditions under which the loss is non-decreasing with the number of rounds, and is strictly decreasing unless the predictor has converged. Second, we show conditions under which the $\alpha$ multicalibration error after $T$ rounds of the algorithm is bounded in terms of the gap between predictors $F_{T+1}$ and $F_{T}$, achieving 0-multicalibration when the predictor has converged.
The results in this section account for the fact that, in each round $t$, the GBM iteratively adds regression trees to $F_{t}$ to form $F_{t+1}$. This is different from the arguments made in Section 3.1 where we assume that an optimal linear combination of weak learners, minimizing the loss function, is added in each round of the algorithm.
We consider a GBM that in round $t$ recursively combines $M_{t}$ regression trees as follows:
$F_{t}^{m+1}(x)=F_{t}^{m}(x)+\rho_{m}h_{j_{m}}(x,f_{t}(x))$
where $j_{m}$ is the index of the regression tree added in the $m$-th iteration, $\rho_{m}$ is the step size, and $m=0,1,\ldots,M_{t}-1$.
Let
$\mathcal{L}_{t}(w)=\mathbb{E}_{D}\left[\mathcal{L}\left(F_{t}(X)+\sum_{k=1}^{K}w_{k}h_{k}(X,f_{t}(X)),Y\right)\right].$
The gradient of the loss function with respect to $w_{k}$ is equal to
$\nabla_{k}\mathcal{L}_{t}(w)=\mathbb{E}_{D}\left[h_{k}(X,f_{t}(X))(Y-f_{t+1}(X))\right]$
where $f_{t+1}(x)=\mu(F_{t}(x)+\sum_{k=1}^{K}w_{k}h_{k}(x,f_{t}(x)))$. A GBM can be shown to correspond to a coordinate gradient descent in the coefficient space (Lu and Mazumder, 2020), defined by
$w^{m+1}=w^{m}-\rho\nabla_{j_{m}}\mathcal{L}_{t}(w^{m})e_{j_{m}}$
where $e_{j}$ is the $j$-th basis vector in $\mathbb{R}^{K}$ and $\rho$ is a step size.
A.2.1. Decreasing Loss Property
Proposition A.0.
Assume that the loss function $\mathcal{L}$ is a convex, $L$-smooth loss function (for log loss, $L=1/4$) and the step size $\rho$ is set as $\rho=1/L$. Then, for every $t\geq 0$,
$\mathbb{E}_{D}\left[\mathcal{L}(F_{t+1}(X),Y)\right]\leq\mathbb{E}_{D}\left[\mathcal{L}(F_{t}(X),Y)\right]$
where the inequality is strict whenever $\nabla_{j_{m}}\mathcal{L}_{t}(w^{m})\neq 0$ in some iteration $m\in\{0,1,\ldots,M_{t}-1\}$.
Proof.
First, note the following relations, for any $\rho_{m}$:
$\mathcal{L}_{t}(w^{m+1})$ $\leq$ $\mathcal{L}_{t}(w^{m}+\rho_{m}e_{j_{m}})$
$\leq$ $\mathcal{L}_{t}(w^{m})+\rho_{m}\nabla\mathcal{L}_{t}(w^{m})^{\top}e_{j_{m}}+\frac{L}{2}\rho_{m}^{2}.$
By taking $\rho_{m}=-(1/L)\nabla_{j_{m}}\mathcal{L}_{t}(w^{m})$, we have
$\mathcal{L}_{t}(w^{m+1})\leq\mathcal{L}_{t}(w^{m})-\frac{1}{2L}(\nabla_{j_{m}}\mathcal{L}_{t}(w^{m}))^{2}.$
It follows that
$\mathcal{L}_{t}(w^{M_{t}})-\mathcal{L}_{t}(w^{0})\leq-\frac{1}{2L}\sum_{m=0}^{M_{t}-1}(\nabla_{j_{m}}\mathcal{L}_{t}(w^{m}))^{2}.$
Hence, it follows that $\mathcal{L}_{t}(w^{M_{t}})-\mathcal{L}_{t}(w^{0})\leq 0$ where the inequality is strict whenever $\nabla_{j_{m}}\mathcal{L}_{t}(w^{m})\neq 0$ for some iteration $m\in\{0,1,\ldots,M_{t}-1\}$. ∎
For a GBM that in each iteration greedily selects a weak learner with the largest absolute value of the loss-gradient, i.e.,
$j_{m}\in\arg\max_{j}|\nabla_{j}\mathcal{L}_{t}(w^{m})|,$
we can establish a rate of convergence to the loss of the optimum linear combination of weak learners, using Theorem 4.2 (Lu and Mazumder, 2020). Let $w^{*}(t)=\arg\min_{w}\mathcal{L}_{t}(w)$. Then,
(5) $\mathbb{E}_{D}\left[\mathcal{L}(F_{t+1}(X),Y)\right]-\mathcal{L}_{t}(w^{*}(t))\leq 0.5\cdot\mathrm{Dist_{t}^{2}}\frac{1}{M_{t}}$
where $\mathrm{Dist}_{t}$ is a distance between the level set $\{w:\mathcal{L}_{t}(w)\leq\mathcal{L}_{t}(w^{0})\}$ and $w^{*}(t)$, given in (Lu and Mazumder, 2020). This provides a theoretical justification for assuming updating the predictor with an optimal ensemble of weak learners, minimizing the loss function, in each round of the algorithm.
A.2.2. A Bound on the Multicalibration Error
We next provide a bound on the multicalibration error of the predictor $f_{T}$.
Proposition A.0.
For any $T>0$, $f_{T}$ is $\alpha$-MC with
$\alpha\leq\frac{2}{\sqrt{3}}C_{T}\sqrt{2\ \mathbb{E}\left[|F_{T+1}(X)-F_{T}(X)|\right]+\frac{1}{8}\mathbb{E}\left[(F_{T+1}(X)-F_{T}(X))^{2}\right]+\epsilon_{T}}$
where
$C_{T}^{2}:=\max_{h\in\mathcal{H},g\in\mathcal{G}}\frac{\mathbb{E}\left[g(X)\mid h(X)=1\right]}{\mathbb{E}\left[f_{T}(X)(1-f_{T}(X))\mid h(X)=1\right]}$
and $\epsilon_{T}:=\mathbb{E}\left[\mathcal{L}(F_{T+1}(X),Y)\right]-\mathcal{L}_{t}(w^{*}(T))$.
Before presenting the proof of the proposition, we discuss the result of the proposition.
The proposition provides a bound on the miscalibration error in terms of the gap between successive predictor values $F_{T+1}$ and $F_{T}$, $\epsilon_{T}$ that captures the gap between the loss function values achieved by the ensemble of regression trees added by the GBM and the optimal ensemble in round $T$, and $C_{T}$ that depends on how much the predictor $f_{T}$ concentrates around the boundary values 0 and $1$. The dependence on the gap between $F_{T}$ and $F_{T+1}$ is such that the multicalibration error decreases as the gap between $F_{T}$ and $F_{T+1}$ decreases. The error term $\epsilon_{T}$ can be controlled by the number of iterations $M_{t}$ used by GBM in round $t$ of the algorithm, as demonstrated in Equation (5).
For the $C_{T}$ term, note:
$C_{T}^{2}\leq\frac{1}{\min_{h\in\mathcal{H}}\mathbb{E}\left[f_{T}(X)(1-f_{T}(X))\mid h(X)=1\right]}.$
This bound is finite if and only if $\mathbb{E}\left[f_{T}(X)(1-f_{T}(X))\mid h(X)=1\right]>0$ for every group $h\in\mathcal{H}$. In other words, for every group $h$, there exists a member $x$ such that $0<f_{T}(x)<1$. This condition clearly holds if $0<f_{T}(x)<1$ for every $x\in\mathcal{X}$ which holds whenever $F_{T}(x)\in[-R,R]$ for every $x\in\mathcal{X}$, for some constant $R>0$. If the predictor $f_{T}$ takes values in $\epsilon,1-\epsilon]$, for some $0<\epsilon\leq 1/2$, then clearly $C_{T}^{2}\leq 2/\epsilon$.
Proof.
We first show the following lemma:
Lemma A.4.
If $f_{T}$ is such that $|\mathbb{E}\left[h(X)g(f_{T}(X))(y-f_{T}(X))\right]|\geq\alpha\sqrt{\mathbb{E}\left[h(X)f_{T}(X)(1-f_{T}(X))\right]}$, for some $h\in\mathcal{H}$ and $g\in\mathcal{G}$, then there exists $\tilde{w}\in\mathbb{R}^{K}$ such that for $\tilde{h}(x)=\sum_{k=1}^{K}\tilde{w}_{k}h_{k}(x,f_{t}(x))$, we have
$\alpha\leq\frac{2}{\sqrt{3}}C_{T}\sqrt{\mathbb{E}\left[\mathcal{L}(F_{T}(X),Y)\right]-\mathbb{E}\left[\mathcal{L}(F_{T}(X)+\tilde{h}(X),Y)\right]}$
where
$C_{T}^{2}:=\max_{h\in\mathcal{H},g\in\mathcal{G}}\frac{\mathbb{E}\left[g(X)\mid h(X)=1\right]}{\mathbb{E}\left[f_{T}(X)(1-f_{T}(X))\mid h(X)=1\right]}.$
Proof.
For any convex, $L$-smooth loss function $\mathcal{L}$, and any $F$ and $h$, we have
$\mathcal{L}(F(x),y)-\mathcal{L}(F(x)+h(x),y)$
$\geq$ $\frac{d}{dF}\mathcal{L}(F(x)+h(x),y)(-h(x))$
$\geq$ $\frac{d}{dF}\mathcal{L}(F(x),y)(-h(x))-Lh(x)^{2}$
where the first inequality holds as $\mathcal{L}$ is convex and the second inequality holds because $\mathcal{L}$ is $L$-smooth.
For the log loss $(d/dF)\mathcal{L}(F,y)=-(y-\sigma(F))$, hence, we have
$\mathbb{E}\left[\mathcal{L}(F(X),Y)\right]-\mathbb{E}\left[\mathcal{L}(F(X)+h(X),Y)\right]$
$\geq$ $\mathbb{E}\left[(Y-f(X))h(X)\right]-L\mathbb{E}\left[h(X)^{2}\right].$
Let $\tilde{w}$ be defined as:
$\tilde{w}_{h,g}=\frac{\mathbf{E}[h(X)g(f_{t}(X))(Y-f_{t}(X))]}{\mathbb{E}\left[(h(X)g(f_{t}(X)))^{2}\right]}$
if $(h,g)=(h^{*},g^{*})$ and $\tilde{\theta}_{h,g}=0$, otherwise, where
$(h^{*},g^{*})\in\arg\max_{h\in\mathcal{G},g\in\mathcal{G}}\frac{|\mathbb{E}\left[h(X)g(f_{t}(X))(Y-f_{t}(X))\right]|}{\sqrt{\mathbb{E}\left[h(X)f_{t}(X)(1-f_{t}(X))\right]}}.$
For $w=\tilde{w}$, we have
$\mathbb{E}\left[(Y-f_{t}(X))\sum_{k=1}^{K}w_{k}h_{k}(x,f_{t}(X))\right]=\frac{\mathbb{E}\left[h^{*}(X)g^{*}(f_{t}(X))(Y-f_{t}(X))\right]^{2}}{\mathbb{E}\left[(h^{*}(X)g^{*}(f_{t}(X)))^{2}\right]}$
and
$\mathbb{E}\left[\left(\sum_{k=1}^{K}w_{k}h_{k}(X,f_{t}(X))\right)^{2}\right]=\frac{\mathbb{E}\left[h^{*}(X)g^{*}(f_{t}(X))(Y-f_{t}(X))\right]^{2}}{\mathbb{E}\left[(h^{*}(X)g^{*}(f_{t}(X)))^{2}\right]}.$
Hence, we have
$\mathbb{E}\left[\mathcal{L}(F_{t}(X),Y)\right]-\mathbb{E}\left[\mathcal{L}\left(F_{t}(X)+\sum_{k=1}^{K}\tilde{w}_{k}h_{k}(X,f_{t}(X)),Y\right)\right]$
$\geq$ $\left(1-L\right)\frac{\mathbb{E}\left[h^{*}(X)g^{*}(f_{t}(X))(Y-f_{t}(X))\right]^{2}}{\mathbb{E}\left[(h^{*}(X)g^{*}(f_{t}(X)))^{2}\right]}$
$\geq$ $\left(1-L\right)\frac{\mathbb{E}\left[h^{*}(X)g^{*}(f_{t}(X))(Y-f_{t}(X))\right]^{2}}{\mathbb{E}\left[h^{*}(X)^{2}\right]}$
$=$ $\left(1-L\right)\frac{\mathbb{E}\left[h^{*}(X)f_{t}(X)(1-f_{t}(X))\right]}{\mathbb{E}\left[h^{*}(X)g^{*}(f_{t}(X))\right]}\cdot$
$\cdot\left(\frac{\mathbb{E}\left[h^{*}(X)g^{*}(f_{t}(X))(Y-f_{t}(X))\right]}{\sqrt{\mathbb{E}\left[h^{*}(X)f_{t}(X)(1-f_{t}(X)\right]}}\right)^{2}$
$\geq$ $\left(1-L\right)\frac{\mathbb{E}\left[f_{t}(X)(1-f_{t}(X))\mid h^{*}(X)=1\right]}{\mathbb{E}\left[g^{*}(f_{t}(X))\mid h^{*}(X)=1\right]}\alpha^{2}.$
Using the fact that for the log loss function $L=1/4$ and the definition of $C_{T}$, we have
$\alpha^{2}$ $\leq$ $\frac{4}{3}C_{T}^{2}\left(\mathbb{E}\left[\mathcal{L}(F_{t}(X),Y)\right]-\mathbb{E}\left[\mathcal{L}\left(F_{t}(X)+\tilde{h}(X),Y\right)\right]\right)$
which completes the proof of the lemma. ∎
For any $w\in\mathbb{R}^{K}$,
$\mathbb{E}\left[\mathcal{L}(F_{t}(X),Y)\right]-\mathbb{E}\left[\mathcal{L}\left(F_{t}(X)+\sum_{k=1}^{K}w_{k}h_{k}(X,f_{t}(X)),Y\right)\right]$
$\leq$ $\mathbb{E}\left[\mathcal{L}(F_{t}(X),Y)\right]-\mathcal{L}_{t}(w^{*}(t))$
$=$ $\mathbb{E}\left[\mathcal{L}(F_{t}(X),Y)\right]-\mathbb{E}\left[\mathcal{L}(F_{t+1}(X),Y)\right]$
$+\mathbb{E}\left[\mathcal{L}(F_{t+1}(X),Y)\right]-\mathcal{L}_{t}(w^{*}(t)).$
We next consider the two terms, $\mathbb{E}\left[\mathcal{L}(F_{t}(X),Y)\right]-\mathbb{E}\left[\mathcal{L}(F_{t+1}(X),Y)\right]$ and $\mathbb{E}\left[\mathcal{L}(F_{t+1}(X),Y)\right]-\mathcal{L}_{t}(w^{*}(t))$. The first term captures the gap of the loss function values at two successive rounds of the algorithm. This term can be bounded as follows:
$\mathbb{E}\left[\mathcal{L}(F_{t}(X),Y)\right]-\mathbb{E}\left[\mathcal{L}(F_{t+1}(X),Y)\right]$
$\leq$ $\mathbb{E}\left[\frac{d}{dF}\mathcal{L}(F_{t+1}(X),Y)\cdot(F_{t}(X)-F_{t+1}(X))+\frac{L}{2}(F_{t}(X)-F_{t+1}(X))^{2}\right]$
$=$ $\mathbb{E}\left[(y-f_{t+1})(F_{t}(X)-F_{t+1}(X))+\frac{L}{2}(F_{t}(X)-F_{t+1}(X))^{2}\right]$
$\leq$ $\mathbb{E}\left[|Y-f_{t+1}(X)||F_{t}(X)-F_{t+1}(X)|+\frac{L}{2}(F_{t}(X)+F_{t+1}(X))^{2}\right]$
$\leq$ $2\ \mathbb{E}\left[|F_{t+1}(X)-F_{t}(X|\right]+\frac{L}{2}\mathbb{E}\left[(F_{t+1}(X)-F_{t}(X))^{2}\right].$
The second term, $\mathbb{E}\left[\mathcal{L}(F_{t+1}(X),Y)\right]-\mathcal{L}_{t}(w^{*}(t))$ is the gap between the loss function value at the end of the $t$-th round of the algorithm and the loss function value obtained by adding an optimal linear combination of weak classifiers. This term can be controlled by the number of iterations used by GBM in round $t$. Specifically, this term can be bounded as given in Equation (5).
The assertion of the proposition follows by combining the above relations with Lemma A.4. ∎
Appendix B Benchmark Datasets
We include in our experimental analysis a total of $11$ datasets. Specifically:
  • American Community Survey (Ding et al., 2021): We use the folktable python package (Flood et al., 2021) to access the data and define six binarized prediction tasks (i.e., create six datasets): income (ACSIncome), mobility (ACSMobility), employment (ACSEmploy), travel time (ACSTravel), public health insurance (ACSPublic), health insurance (ACSHealth) . Also, ACSIncomeCA considers only the rows with the country set to California. For each target we use the features specified in the folktable package and add all features used in the specifications protected groups in Hansen et al. (2024);
  • UCI Bank Marketing: The dataset Marketing has the task of predicting if a client will subscribe to a term deposit, using features like housing, occupation, education, and age (Moro et al., 2014);
  • UCI Default of Credit Card Clients: The dataset Credit has the task of predicting credit card debt default using payment history and identity attributes (Yeh, 2009);
  • Home Mortgage Disclosure Act: The dataset HMDA has the task of predicting the acceptance of mortgage applications in 2017, using identity attributes of associated applicants (Cooper et al., 2023);
  • Medical Expenditure Panel Survey: The dataset MEPS has the task of predicting whether a household makes at least $10$ medical visits, based on socioeconomic and geographic information (Sharma et al., 2021).
Table 4 shows the number of samples, features and the class prior of each dataset.
Dataset # Samples $n$ # Features $d$ Class Distr.
MEPS 11079 139 0.169
Credit 30000 118 0.221
Marketing 45211 41 0.116
HMDA 114185 89 0.752
ACSIncomeCA 195665 10 0.410
ACSMobility 616207 23 0.735
ACSPublic 1123374 24 0.293
ACSTravel 1458542 19 0.437
ACSIncome 1655429 10 0.370
ACSEmploy 3207990 20 0.456
ACSHealth 3207990 31 0.150
Table 4. Number of samples, features and class distribution ($p(Y=1)$) for each dataset used in the benchmark evaluation.
Appendix C Setting Default Hyperparameters for MCGrad
MCGrad uses LightGBM as GBDT implementation. We found that order to get ML engineers to adopt MCGrad into their ML production system, is important that it works out-of-the-box. That is, it should work for their ML model, without application-specific tuning of the LightGBM hyperparameters. Making it possible for ML engineers to use MCGrad without tuning LightGBM hyperparameters substantially reduces the barrier to use MCGrad, and increases trust that it can be safely incorporated into automated ML model training pipelines.
In order to set default hyperparameters for MCGrad, we performed a grid search of LightGBM hyperparameters where we evaluated each hyperparameter configuration on a collection of 35 company-internal datasets. The purpose of this analysis was to select a set of LightGBM parameters that are suitable as default hyperparameters in MCGrad. We selected the hyperparameter configuration that across those 35 datasets (1) never degraded log loss and Area Under the Precision-Recall Curve (AUPRC) relative to the model scores before calibration adjustment, and (2) minimised the average MultiCalibration Error (MCE) across the datasets.
This resulted in the following set of hyperparameters:
learning_rate:
$0.02873$
max_depth:
55
min_child_samples:
$160$
n_estimators:
$94$
num_leaves:
55
lambda_l2:
$0.00913$
min_gain_to_split:
$0.15$
All other hyperparameters were left to LightGBM default values.
dataset mcb_algorithm $\text{MCE}*\sigma$ (unspecified) MCE (unspecified) $\text{MCE}*\sigma$ (prespecified) MCE (prespecified) max smECE (prespecified)
MEPS base_model 0.0265 3.9940 0.0926 3.2849 0.0819
MCGrad_msh_20 0.0253 3.8815 0.0815 3.2923 0.0755
MCGrad 0.0265 3.9940 0.0926 3.2849 0.0819
MCGrad_group_features 0.0265 3.9940 0.0926 3.2849 0.0819
MCGrad_no_unshrink 0.0265 3.9940 0.0926 3.2849 0.0819
MCGrad_one_round 0.0256 3.9250 0.0806 3.2888 0.0733
DFMC 0.0266 4.0473 0.0870 3.4147 0.0761
HKRR 0.0318 4.8624 0.0872 3.1814 0.1054
Isotonic 0.0237 3.6076 0.0797 3.4070 0.0713
acs_employment_all_states base_model 0.0274 87.2951 0.0263 45.5082 0.0499
MCGrad_msh_20 0.0019 7.1680 0.0099 4.7747 0.0137
MCGrad 0.0085 30.7752 0.0126 14.2049 0.0199
MCGrad_group_features 0.0143 47.9972 0.0061 3.3573 0.0120
MCGrad_no_unshrink 0.0085 31.1755 0.0128 14.3882 0.0201
MCGrad_one_round 0.0085 30.7752 0.0126 14.2049 0.0199
DFMC 0.0156 52.0332 0.0110 12.6897 0.0152
HKRR 0.0196 63.8249 0.0104 14.3973 0.0110
Isotonic 0.0153 50.7122 0.0210 30.8645 0.0215
HMDA base_model 0.0283 13.9421 0.0475 5.5320 0.0468
MCGrad_msh_20 0.0091 4.7703 0.0300 3.2015 0.0365
MCGrad 0.0093 4.8958 0.0309 3.2932 0.0371
MCGrad_group_features 0.0283 13.9249 0.0372 2.8149 0.0363
MCGrad_no_unshrink 0.0080 4.1592 0.0325 3.0728 0.0405
MCGrad_one_round 0.0096 4.8650 0.0443 3.0720 0.0504
DFMC 0.0287 14.0359 0.0372 2.8040 0.0380
HKRR 0.0285 13.8664 0.0401 3.3343 0.0380
Isotonic 0.0301 14.7130 0.0468 4.6516 0.0456
acs_travel_time_all_states base_model 0.0192 21.5562 0.1736 18.6954 0.1688
MCGrad_msh_20 0.0059 6.8154 0.0330 4.9704 0.0346
MCGrad 0.0059 6.8237 0.0331 4.9730 0.0346
MCGrad_group_features 0.0165 18.6758 0.0297 2.7626 0.0293
MCGrad_no_unshrink 0.0061 6.9743 0.0349 4.1746 0.0365
MCGrad_one_round 0.0118 13.3776 0.0865 12.9763 0.0840
DFMC 0.0166 18.7304 0.0331 4.2345 0.0323
HKRR 0.0164 18.4043 0.0355 12.7941 0.0349
Isotonic 0.0167 18.7715 0.2116 18.1106 0.2025
acs_income_all_states base_model 0.0233 34.0505 0.0714 17.9622 0.0708
MCGrad_msh_20 0.0049 7.9316 0.0318 6.2694 0.0314
MCGrad 0.0050 8.1133 0.0322 6.5280 0.0317
MCGrad_group_features 0.0201 29.6245 0.0106 5.3609 0.0152
MCGrad_no_unshrink 0.0049 7.8965 0.0317 4.5370 0.0305
MCGrad_one_round 0.0179 27.7918 0.0520 13.9878 0.0502
DFMC 0.0211 31.0460 0.0185 9.9542 0.0204
HKRR 0.0230 33.8068 0.0408 10.2374 0.0413
Isotonic 0.0216 31.7564 0.0701 17.3586 0.0694
ACSIncome base_model 0.0265 13.4846 0.0611 6.1578 0.0579
MCGrad_msh_20 0.0104 6.0717 0.0443 3.8099 0.0498
MCGrad 0.0102 5.9816 0.0439 3.8675 0.0501
MCGrad_group_features 0.0239 12.3216 0.0320 3.4824 0.0337
MCGrad_no_unshrink 0.0109 6.2974 0.0446 3.7393 0.0530
MCGrad_one_round 0.0187 10.3174 0.0489 5.6888 0.0520
DFMC 0.0243 12.5025 0.0345 3.7487 0.0351
HKRR 0.0246 12.5983 0.0338 3.6878 0.0354
Isotonic 0.0245 12.6166 0.0592 5.1895 0.0574
acs_health_insurance_all_states base_model 0.0198 45.2690 0.0803 38.2892 0.0800
MCGrad_msh_20 0.0044 10.4370 0.0082 4.0131 0.0108
MCGrad 0.0044 10.4370 0.0082 4.0131 0.0108
MCGrad_group_features 0.0120 27.8627 0.0038 3.2136 0.0075
MCGrad_no_unshrink 0.0045 10.5774 0.0093 3.8013 0.0104
MCGrad_one_round 0.0096 22.4316 0.0522 21.3152 0.0520
DFMC 0.0118 27.4016 0.0148 7.5026 0.0232
HKRR 0.0150 33.9543 0.0322 17.3882 0.0319
Isotonic 0.0134 30.9506 0.0780 31.8857 0.0778
acs_public_health_insurance_all_states base_model 0.0354 40.4553 0.1035 22.1733 0.1025
MCGrad_msh_20 0.0068 8.3577 0.0324 4.9253 0.0400
MCGrad 0.0067 8.2452 0.0320 4.9934 0.0396
MCGrad_group_features 0.0282 32.7613 0.0305 3.4484 0.0324
MCGrad_no_unshrink 0.0076 9.2211 0.0289 5.3953 0.0380
MCGrad_one_round 0.0265 31.3918 0.0668 12.7760 0.0654
DFMC 0.0277 32.0694 0.0411 6.4715 0.0434
HKRR 0.0342 39.5956 0.0428 7.3886 0.0403
Isotonic 0.0329 38.0286 0.1064 21.8710 0.1051
acs_mobility_all_states base_model 0.0642 52.3930 0.0689 9.0919 0.0652
MCGrad_msh_20 0.0068 5.8904 0.0306 3.1756 0.0349
MCGrad 0.0068 5.8824 0.0302 3.2353 0.0347
MCGrad_group_features 0.0673 55.1968 0.0156 2.4286 0.0200
MCGrad_no_unshrink 0.0060 5.2134 0.0338 2.9911 0.0370
MCGrad_one_round 0.0150 12.7685 0.0423 6.9712 0.0414
DFMC 0.0671 54.9004 0.0217 2.9196 0.0261
HKRR 0.0670 54.5418 0.0406 5.2062 0.0433
Isotonic 0.0669 54.7596 0.0659 5.2414 0.0677
CreditDefault base_model 0.0168 3.5652 0.0783 2.1844 0.0536
MCGrad_msh_20 0.0162 3.4353 0.0842 2.2523 0.0543
MCGrad 0.0162 3.4467 0.0790 2.1717 0.0536
MCGrad_group_features 0.0168 3.5723 0.0772 2.3468 0.0526
MCGrad_no_unshrink 0.0164 3.4813 0.0796 2.2148 0.0536
MCGrad_one_round 0.0162 3.4467 0.0790 2.1717 0.0536
DFMC 0.0164 3.4734 0.0943 2.2036 0.0596
HKRR 0.0180 3.7981 0.0559 2.5684 0.0538
Isotonic 0.0168 3.5659 0.0847 2.3044 0.0544
BankMarketing base_model 0.0195 7.0101 0.0522 5.1177 0.0565
MCGrad_msh_20 0.0156 5.9140 0.0239 1.8957 0.0328
MCGrad 0.0110 4.1535 0.0257 1.8693 0.0321
MCGrad_group_features 0.0163 5.8165 0.0256 2.1236 0.0362
MCGrad_no_unshrink 0.0103 3.8441 0.0222 2.1485 0.0317
MCGrad_one_round 0.0139 5.0625 0.0384 3.0142 0.0415
DFMC 0.0158 5.6785 0.0267 1.9211 0.0350
HKRR 0.0270 9.1271 0.0360 3.9728 0.0390
Isotonic 0.0177 6.3438 0.0288 2.4565 0.0370
Table 5. Multicalibration metrics.
dataset mcb_algorithm ECCE $\text{ECCE}/\sigma$ ECE Log loss PRAUC
MEPS base_model 0.0157 2.3774 0.0215 0.3177 0.6211
MCGrad_msh_20 0.0146 2.2479 0.0186 0.3150 0.6221
MCGrad 0.0157 2.3774 0.0215 0.3177 0.6211
MCGrad_group_features 0.0157 2.3774 0.0215 0.3177 0.6211
MCGrad_no_unshrink 0.0157 2.3774 0.0215 0.3177 0.6211
MCGrad_one_round 0.0140 2.1413 0.0189 0.3153 0.6204
DFMC 0.0143 2.1821 0.0160 0.3163 0.6116
HKRR 0.0166 2.5308 0.0252 0.3367 0.5317
Isotonic 0.0112 1.7140 0.0118 0.3160 0.5945
acs_employment_all_states base_model 0.0162 51.6562 0.0328 0.2257 0.9273
MCGrad_msh_20 0.0010 3.8704 0.0013 0.1511 0.9597
MCGrad 0.0012 4.4955 0.0023 0.1599 0.9512
MCGrad_group_features 0.0008 2.5486 0.0009 0.2001 0.9391
MCGrad_no_unshrink 0.0013 4.7459 0.0027 0.1588 0.9519
MCGrad_one_round 0.0038 13.7094 0.0067 0.1698 0.9451
DFMC 0.0011 3.7251 0.0025 0.2181 0.9349
HKRR 0.0031 10.1073 0.0062 0.2111 0.9238
Isotonic 0.0009 2.9845 0.0011 0.2066 0.9294
HMDA base_model 0.0036 1.7643 0.0063 0.2761 0.9772
MCGrad_msh_20 0.0044 2.3081 0.0057 0.2539 0.9818
MCGrad 0.0048 2.5294 0.0060 0.2541 0.9818
MCGrad_group_features 0.0029 1.4001 0.0030 0.2746 0.9775
MCGrad_no_unshrink 0.0034 1.7673 0.0039 0.2546 0.9817
MCGrad_one_round 0.0033 1.6494 0.0055 0.2609 0.9805
DFMC 0.0032 1.5643 0.0055 0.2747 0.9775
HKRR 0.0047 2.3059 0.0069 0.2773 0.9723
Isotonic 0.0023 1.1215 0.0024 0.2757 0.9760
acs_travel_time_all_states base_model 0.0093 10.4042 0.0181 0.6520 0.5258
MCGrad_msh_20 0.0028 3.2589 0.0053 0.6304 0.5775
MCGrad 0.0028 3.2067 0.0053 0.6303 0.5776
MCGrad_group_features 0.0013 1.4963 0.0015 0.6449 0.5337
MCGrad_no_unshrink 0.0021 2.4608 0.0037 0.6294 0.5781
MCGrad_one_round 0.0053 5.9904 0.0068 0.6381 0.5579
DFMC 0.0025 2.8485 0.0047 0.6460 0.5318
HKRR 0.0051 5.7510 0.0052 0.6480 0.5083
Isotonic 0.0032 3.5899 0.0007 0.6487 0.5233
acs_income_all_states base_model 0.0067 9.7652 0.0135 0.4706 0.7407
MCGrad_msh_20 0.0019 3.0261 0.0029 0.3924 0.8299
MCGrad 0.0020 3.1717 0.0029 0.3924 0.8298
MCGrad_group_features 0.0013 1.9035 0.0014 0.4608 0.7503
MCGrad_no_unshrink 0.0012 1.9568 0.0018 0.3926 0.8296
MCGrad_one_round 0.0034 5.2775 0.0049 0.4230 0.7963
DFMC 0.0019 2.7820 0.0029 0.4648 0.7469
HKRR 0.0032 4.7310 0.0038 0.4681 0.7230
Isotonic 0.0015 2.1644 0.0018 0.4674 0.7402
ACSIncome base_model 0.0076 3.8599 0.0129 0.4635 0.7898
MCGrad_msh_20 0.0064 3.7358 0.0120 0.3863 0.8599
MCGrad 0.0065 3.8167 0.0114 0.3862 0.8600
MCGrad_group_features 0.0019 1.0030 0.0046 0.4556 0.7963
MCGrad_no_unshrink 0.0041 2.3710 0.0086 0.3853 0.8601
MCGrad_one_round 0.0035 1.9324 0.0080 0.4076 0.8398
DFMC 0.0019 0.9585 0.0045 0.4568 0.7955
HKRR 0.0023 1.1808 0.0070 0.4604 0.7721
Isotonic 0.0024 1.2285 0.0052 0.4605 0.7872
acs_health_insurance_all_states base_model 0.0091 20.7777 0.0172 0.3971 0.2730
MCGrad_msh_20 0.0011 2.5741 0.0019 0.3762 0.3331
MCGrad 0.0011 2.5656 0.0018 0.3762 0.3334
MCGrad_group_features 0.0005 1.1681 0.0008 0.3898 0.2933
MCGrad_no_unshrink 0.0007 1.6650 0.0015 0.3762 0.3330
MCGrad_one_round 0.0054 12.5842 0.0067 0.3844 0.3072
DFMC 0.0018 4.1372 0.0055 0.3909 0.2912
HKRR 0.0073 16.6308 0.0069 0.3944 0.2567
Isotonic 0.0009 2.0126 0.0007 0.3951 0.2702
acs_public_health_insurance_all_states base_model 0.0160 18.1469 0.0270 0.5194 0.5979
MCGrad_msh_20 0.0013 1.6389 0.0022 0.4682 0.6684
MCGrad 0.0014 1.7647 0.0021 0.4681 0.6686
MCGrad_group_features 0.0018 2.0503 0.0024 0.5082 0.6091
MCGrad_no_unshrink 0.0020 2.3991 0.0023 0.4672 0.6703
MCGrad_one_round 0.0097 11.4243 0.0105 0.4934 0.6328
DFMC 0.0024 2.7898 0.0029 0.5101 0.6056
HKRR 0.0054 6.2067 0.0105 0.5147 0.5595
Isotonic 0.0019 2.2277 0.0012 0.5144 0.5921
acs_mobility_all_states base_model 0.0076 6.2231 0.0173 0.5540 0.8059
MCGrad_msh_20 0.0054 4.6724 0.0080 0.5092 0.8694
MCGrad 0.0053 4.6375 0.0079 0.5092 0.8695
MCGrad_group_features 0.0021 1.6867 0.0036 0.5507 0.8112
MCGrad_no_unshrink 0.0032 2.7402 0.0050 0.5092 0.8691
MCGrad_one_round 0.0097 8.2706 0.0119 0.5176 0.8612
DFMC 0.0029 2.4097 0.0035 0.5514 0.8095
HKRR 0.0048 3.9479 0.0061 0.5532 0.7936
Isotonic 0.0019 1.5359 0.0016 0.5521 0.8045
CreditDefault base_model 0.0077 1.6218 0.0092 0.4333 0.5319
MCGrad_msh_20 0.0056 1.1842 0.0103 0.4325 0.5320
MCGrad 0.0056 1.1911 0.0110 0.4328 0.5318
MCGrad_group_features 0.0058 1.2361 0.0118 0.4337 0.5291
MCGrad_no_unshrink 0.0051 1.0907 0.0100 0.4327 0.5318
MCGrad_one_round 0.0056 1.1911 0.0110 0.4328 0.5318
DFMC 0.0059 1.2568 0.0129 0.4343 0.5275
HKRR 0.0073 1.5338 0.0134 0.4402 0.4808
Isotonic 0.0065 1.3702 0.0091 0.4445 0.5251
BankMarketing base_model 0.0167 5.9821 0.0320 0.2433 0.5587
MCGrad_msh_20 0.0034 1.2817 0.0062 0.2141 0.6190
MCGrad 0.0031 1.1650 0.0051 0.2133 0.6209
MCGrad_group_features 0.0035 1.2616 0.0039 0.2307 0.5637
MCGrad_no_unshrink 0.0037 1.3911 0.0053 0.2137 0.6167
MCGrad_one_round 0.0073 2.6721 0.0126 0.2255 0.5806
DFMC 0.0035 1.2494 0.0059 0.2303 0.5634
HKRR 0.0102 3.4546 0.0174 0.2453 0.4933
Isotonic 0.0030 1.0791 0.0060 0.2312 0.5491
Table 6. (Global) Calibration and model performance metrics.